# Adaptive Designs in Fast-Track Registration Processes

### v1

@@ Appendix A Proofs
Let $A:\mathbb{R}\rightarrow[0,0.5]$ be a conditional error function which is positive on $[z_{f},\infty)$, $z_{f}>0$, and $\delta=\xi\delta_{\text{rel}}$ for $\xi\geq 0$ be the a priori assumed or true parameter and $I_{1}>0$, $\alpha,\beta\in(0,0.5)$. By setting $A$ constant to $\alpha$, the non-adaptive design can be considered. We define the sucess probability at $\delta$ by
$\mathcal{S}_{A}(I_{2,\min},I_{1})$ $=\mathbb{P}_{\delta}\left(Z_{1}\geq z_{f},\text{reject\leavevmode\nobreak\ }H_{0}\right)$
@@
$I_{2}^{A}(z_{1}):=\max\left\{I_{2,\min},I_{1}{\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-A(z_{1}))\right\}^{2}}/{z_{f}^{2}}\right\},\leavevmode\nobreak\ z_{1}\geq z_{f}\leavevmode\nobreak\ .$
@@
gives a pointwise non-decreasing sequence of functions which converge pointwise to the constant function $z_{1}\mapsto 1$. Then, due to the monotone convergence theorem, $\mathcal{S}_{A}(I_{2,\min},I_{1})\nearrow 1\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})>p$ which implies for sufficiently large $I_{2,\min}$ it holds $\mathcal{S}_{A}(I_{2,\min},I_{1})>p$. It remains to show that there exists a unique minimal information such that equality holds. On the one hand, $\mathcal{S}_{A}(I_{2,\min},I_{1})$ strictly increases in $I_{2,\min}$. This follows from the fact that the used sample size formula for z-tests is monotonically decreasing in $z_{1}$ and converges to zero as $z_{1}$ approaches infinity. Indeed, this yields if $I_{2,\min}$ increases, $I_{2}^{A}$ increases strictly on some non-empty interval $(z_{i},\infty)$, $z_{i}\geq z_{f}$ and thus $\mathcal{S}_{A}(I_{2,\min},I_{1})$ strictly increases. Furthermore, the dominated convergence theorem gives the continuity of $\mathcal{S}_{A}(I_{2,\min},I_{1})$ in $I_{2,\min}$. By putting al that together, we obtain: If $\mathcal{S}_{A}(0,I_{1})\leq p$, we find a unique $I_{2,\min}\geq 0$ such that $\mathcal{S}_{A}(I_{2,\min},I_{1})=p$. ∎
In the main article, we use the previous lemma twice. In the first part of the paper, we consider all $\delta$ and $I_{1}$ which fulfil $\mathbb{P}_{\delta}(Z_{1}\geq z_{f})>1-\beta$ and thus the lemma can be used for $p:=1-\beta$. In the second part of the paper we use that $\mathbb{P}_{\delta}(\text{reject\leavevmode\nobreak\ }H_{0}\leavevmode\nobreak\ |\leavevmode\nobreak\ Z_{1}\geq z_{f})=1-\beta$ is equivalent to $\mathbb{P}_{\delta}(\text{reject\leavevmode\nobreak\ }H_{0},Z_{1}\geq z_{f})=(1-\beta)\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})$ and set $p:=(1-\beta)\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})$.
@@
$\max_{z_{1}\geq z_{f}}I_{2}^{A}(z_{1},I_{2,\min}^{A},I_{1})=\max\left\{I_{2,\min}^{A},I_{1}{\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-A(z_{f}))\right\}^{2}}/{z_{f}^{2}}\right\}\leavevmode\nobreak\ .$ (18)
@@
$I_{2,\min}^{A}\geq\max\nolimits_{z_{1}\geq z_{f}}I_{2}^{\alpha}(z_{1})=\max\left\{I_{2,\min}^{\alpha},{I_{1}\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-\alpha)\right\}^{2}}/{z_{f}^{2}}\right\}\leavevmode\nobreak\ .$ (20)
If $I_{2,\min}^{A}=0$, (20) cannot be true because the second part of the maximum in (20) is greater than 00 due to our assumptions. Then, (17) is proven. If $I_{2,\min}^{A}>0$ the overall power at $\delta$ is exactly $1-\beta$, i.e. $\mathcal{S}_{A}(I_{2,\min}^{A},I_{1})=1-\beta$. If (20) were true, it would imply because (19) that for all $z_{1}\geq z_{f}$ it holds $I_{2}^{A}(z_{1})=I_{2,\min}^{A}$. Due to $A(z_{f})>\alpha$ and due to the monotonicity of $A$ on $[z_{f},\infty)$ which implies $A(z_{1})\geq A(z_{f})>\alpha$ it holds $\Phi^{-1}(1-A(z_{1}))<\Phi^{-1}(1-\alpha)$. Together with the fact that we have assumed (20) and $I_{2}^{\alpha}(z_{1})$ is decreasing in $z_{1}$, it can be concluded that the integrated of $\mathcal{S}_{A}(I_{2,\min}^{A},I_{1})$ in (4.2) is strictly greater than the integrand of $\mathcal{S}(I_{2,\min}^{\alpha},I_{1})$ in (4.1). Therefore, $\mathcal{S}_{A}(I_{2,\min}^{A},I_{1})>\mathcal{S}(I_{2,\min}^{\alpha},I_{1})$. Because $\mathcal{S}_{A}(I_{2,\min}^{A},I_{1})=1-\beta$ and $\mathcal{S}(I_{2,\min}^{\alpha},I_{1})\geq 1-\beta$, we have a contradiction. Thus, (20) does not hold true, and (17) is fulfilled. ∎
The following example refers to Lemma 2.
Example.
Let $\alpha_{c}=0.15$, $\alpha=0.025$, $\beta=0.2$ like in the main article. For the inverse normal method it holds for all $\xi>\xi_{\min}$ and all $I_{1}>I_{1,\min}$ (notice that $z_{f}$ depends on $I_{1}$) that $A_{I,z_{f}}(z_{f})\geq A_{I,-\infty}(z_{f,\min})$ where $z_{f,\min}:=\max\{\sqrt{t_{\xi}(I_{1,\min})}\eta_{f}/\xi,\Phi^{-1}(1-\alpha_{c})\}$ (analogously for $A_{F,z_{f}}$). It can be calculated that $c_{I}(-\infty)=0.0253$, $c_{F}(-\infty)=0.0044$ and it holds $A_{I,-\infty}(z)>\alpha$ iff $z>0.8041$ and $A_{F,-\infty}(z)>\alpha$ iff $z>0.9382$. For $\xi=1.75$ and $\xi=2$ we have $z_{f,\min}=1.1222$ and $z_{f,\min}=1.0364$, respectively. This means for both considered $\xi$ for both adaptive designs for all $I_{1}>I_{1,\min}$ definitely holds $A_{I,z_{f}}(z_{f})>\alpha$ and $A_{F,z_{f}}(z_{f})>\alpha$. Therefore, due to Lemma 2, both adaptive designs have a lower maximum sum of information than the non-adaptive design.
@@ Appendix B Numerical Implementation
Let $I_{2,1}<I_{2,2}$. We prove $\mathbb{P}_{\delta}(Z_{1}<z_{f},Z(I_{2,1})\geq z_{\alpha})<\mathbb{P}_{\delta}(Z_{1}<z_{f},Z(I_{2,2})\geq z_{\alpha})$ by using Slepian’s inequality (see, for instance, [39, Theorem 5.1.7.]). Slepian’s inequality compares the strength of the dependence of two multivariate normally distributed random vectors with the same mean and the same variance and further assumptions on the covariances. We define $X_{1}:=Y_{1}:=Z_{1}-\delta\sqrt{I_{1}}$ and $X_{2}:=-Z(I_{2,1})+\delta\sqrt{I_{1}+I_{2,1}}$, $Y_{2}:=-Z(I_{2,2})+\delta\sqrt{I_{1}+I_{2,2}}$. Then $X=(X_{1},X_{2})$ and $Y=(Y_{1},Y_{2})$ are centered random vectors and all variances equal $1$. Furthermore, $\mathrm{Cov}(X_{1},X_{2})=-\mathrm{Cov}(Z_{1},Z(I_{2,1}))=-\sqrt{I_{1}/{(I_{1}+I_{2,1})}}\leq-\sqrt{I_{1}/{(I_{1}+I_{2,2})}}=-\mathrm{Cov}(Z_{1},Z(I_{2,2}))=\mathrm{Cov}(Y_{1},Y_{2})$. Slepian’s inequality now gives for all $a_{1},a_{2}\in\mathbb{R}$ that $\mathbb{P}_{\delta}(X_{1}<a_{1},X_{2}\leq a_{2})\leq\mathbb{P}_{\delta}(Y_{1}<a_{1},Y_{2}\leq a_{2})$. By setting $a_{1}=z_{f}-\delta\sqrt{I_{1}}$ and $a_{2}=-z_{\alpha}+\delta\sqrt{I_{1}+I_{2,1}}$ and using $\delta>0$ which implies $a_{2}<-z_{\alpha}+\delta\sqrt{I_{1}+I_{2,2}}$, we obtain
@@
$<\mathbb{P}_{\delta}(Y_{1}<a_{1},Y_{2}\leq-z_{\alpha}+\delta\sqrt{I_{1}+I_{2,2}})=\mathbb{P}_{\delta}(Z_{1}<z_{f},Z(I_{2,2})\geq z_{\alpha})\leavevmode\nobreak\ .$
@@ C.2 Without binding futility stop
In this Section, it is assumed as in Section 4 that at the time of the planning of a fast-track procedure by the applicants, a successful conditional registration is a requirement for permanent registration. Thus, we assume that $\mathbb{P}_{\delta}(Z_{1}\geq z_{f},\text{\leavevmode\nobreak\ reject\leavevmode\nobreak\ }H_{0})=1-\beta$ is striven at an a priori assumed or true effect $\delta=\xi\cdot\delta_{\text{rel}}$ where $\xi>\xi_{\min}$. We assume that the designs and informations are chosen such that this success probability is achieved, as in Section 4. In contrast to Section 4, the conditional error functions are not chosen in a way that there is a binding futility stop after unsuccessful conditional registration. Instead, in this section, we deal with this case like an unbinding futility stop and consider $A_{I,-\infty}$ and $A_{F,-\infty}$. These are somewhat conservative designs with somewhat larger sample sizes compared to $A_{I,z_{f}}$ and $A_{F,z_{f}}$. However, such designs provide the opportunity to continue with a second stage for a permanent registration whenever there are good reasons for and sufficient resources have been achieved.
@@
Figure 12: Minimum needed information for the second stage to achieve overall power of $0.81-\beta=$ 0.8 for the common non-adaptive design and the adaptive designs with conditional error functions $A_{I,-\infty}$ (inverse normal method) and $A_{F,-\infty}$ (Fisher’s product test).
@@ Appendix D Additional Plots for Section 5
The plots in this section complete the discussion in Section 5. Two further plots for the effect assumption $\xi=1.25$ are presented. Additionally, all efficiency parameters considered in Section 5 for $\xi=1.25$ are calculated under the assumption $\xi=1.43=\xi_{\min}$.

### v3

@@ Appendix A Proofs
Let $A:\mathbb{R}\rightarrow[0,0.5]$ be a conditional error function which is positive on $z_{f},\infty)$, $z_{f}>0$, and $\delta=\xi\delta_{\text{rel}}$ for $\xi\geq 0$ be the a priori assumed or true parameter and $I_{1}>0$, $\alpha,\beta\in(0,0.5)$. By setting $A$ constant to $\alpha$, the non-adaptive design can be considered. We define the success probability at $\delta$ by
$\mathcal{S}_{A}(I_{2,\min},I_{1})$ $=\mathbb{P}_{\delta}\left(Z_{1}\geq z_{f},\text{reject\penalty 10000\ }H_{0}\right)$
@@
$I_{2}^{A}(z_{1}):=\max\left\{I_{2,\min},I_{1}{\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-A(z_{1}))\right\}^{2}}/{z_{f}^{2}}\right\},\penalty 10000\ z_{1}\geq z_{f}\penalty 10000\ .$
@@
gives a pointwise non-decreasing sequence of functions which converge pointwise to the constant function $z_{1}\mapsto 1$. Then, due to the monotone convergence theorem, $\mathcal{S}_{A}(I_{2,\min},I_{1})\nearrow 1\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})>p$ which implies for sufficiently large $I_{2,\min}$ it holds $\mathcal{S}_{A}(I_{2,\min},I_{1})>p$. It remains to show that there exists a unique minimal information such that equality holds. On the one hand, $\mathcal{S}_{A}(I_{2,\min},I_{1})$ strictly increases in $I_{2,\min}$. This follows from the fact that the used sample size formula for z-tests is monotonically decreasing in $z_{1}$ and converges to zero as $z_{1}$ approaches infinity. Indeed, this yields if $I_{2,\min}$ increases, $I_{2}^{A}$ increases strictly on some non-empty interval $(z_{i},\infty)$, $z_{i}\geq z_{f}$ and thus $\mathcal{S}_{A}(I_{2,\min},I_{1})$ strictly increases. Furthermore, the dominated convergence theorem gives the continuity of $\mathcal{S}_{A}(I_{2,\min},I_{1})$ in $I_{2,\min}$. By putting all that together, we obtain: If $\mathcal{S}_{A}(0,I_{1})\leq p$, we find a unique $I_{2,\min}\geq 0$ such that $\mathcal{S}_{A}(I_{2,\min},I_{1})=p$. ∎
In the main article, we use the previous lemma twice. In the first part of the paper, we consider all $\delta$ and $I_{1}$ which fulfill $\mathbb{P}_{\delta}(Z_{1}\geq z_{f})>1-\beta$ and thus the lemma can be used for $p:=1-\beta$. In the second part of the paper we use that $\mathbb{P}_{\delta}(\text{reject\penalty 10000\ }H_{0}\penalty 10000\ |\penalty 10000\ Z_{1}\geq z_{f})=1-\beta$ is equivalent to $\mathbb{P}_{\delta}(\text{reject\penalty 10000\ }H_{0},Z_{1}\geq z_{f})=(1-\beta)\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})$ and set $p:=(1-\beta)\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})$.
@@
$\max_{z_{1}\geq z_{f}}I_{2}^{A}(z_{1},I_{2,\min}^{A},I_{1})=\max\left\{I_{2,\min}^{A},I_{1}{\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-A(z_{f}))\right\}^{2}}/{z_{f}^{2}}\right\}\penalty 10000\ .$ (19)
@@
$I_{2,\min}^{A}\geq\max\nolimits_{z_{1}\geq z_{f}}I_{2}^{\alpha}(z_{1})=\max\left\{I_{2,\min}^{\alpha},{I_{1}\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-\alpha)\right\}^{2}}/{z_{f}^{2}}\right\}\penalty 10000\ .$ (21)
If $I_{2,\min}^{A}=0$, (21) cannot be true because the second part of the maximum in (21) is greater than 0 due to our assumptions. Then, (18) is proven. If $I_{2,\min}^{A}>0$ the overall power at $\delta$ is exactly $1-\beta$, i.e. $\mathcal{S}_{A}(I_{2,\min}^{A},I_{1})=1-\beta$. If (21) were true, it would imply because (20) that for all $z_{1}\geq z_{f}$ it holds $I_{2}^{A}(z_{1})=I_{2,\min}^{A}$. Due to $A(z_{f})>\alpha$ and due to the monotonicity of $A$ on $z_{f},\infty)$ which implies $A(z_{1})\geq A(z_{f})>\alpha$ it holds $\Phi^{-1}(1-A(z_{1}))<\Phi^{-1}(1-\alpha)$. Together with the fact that we have assumed (21) and $I_{2}^{\alpha}(z_{1})$ is decreasing in $z_{1}$, it can be concluded that the integrated of $\mathcal{S}_{A}(I_{2,\min}^{A},I_{1})$ in (4.2) is strictly greater than the integrand of $\mathcal{S}(I_{2,\min}^{\alpha},I_{1})$ in (4.1). Therefore, $\mathcal{S}_{A}(I_{2,\min}^{A},I_{1})>\mathcal{S}(I_{2,\min}^{\alpha},I_{1})$. Because $\mathcal{S}_{A}(I_{2,\min}^{A},I_{1})=1-\beta$ and $\mathcal{S}(I_{2,\min}^{\alpha},I_{1})\geq 1-\beta$, we have a contradiction. Thus, (21) does not hold true, and (18) is fulfilled. ∎
@@ Appendix B Numerical Implementation
Let $I_{2,1}<I_{2,2}$. We prove $\mathbb{P}_{\delta}(Z_{1}<z_{f},Z(I_{2,1})\geq z_{\alpha})<\mathbb{P}_{\delta}(Z_{1}<z_{f},Z(I_{2,2})\geq z_{\alpha})$ by using Slepian’s inequality (see, for instance, Theorem 5.1.7 in the work of Tong [45]). Slepian’s inequality compares the strength of the dependence of two multivariate normally distributed random vectors with the same mean and the same variance and further assumptions on the covariances. We define $X_{1}:=Y_{1}:=Z_{1}-\delta\sqrt{I_{1}}$ and $X_{2}:=-Z(I_{2,1})+\delta\sqrt{I_{1}+I_{2,1}}$, $Y_{2}:=-Z(I_{2,2})+\delta\sqrt{I_{1}+I_{2,2}}$. Then $X=(X_{1},X_{2})$ and $Y=(Y_{1},Y_{2})$ are centered random vectors and all variances equal $1$. Furthermore, $\mathrm{Cov}(X_{1},X_{2})=-\mathrm{Cov}(Z_{1},Z(I_{2,1}))=-\sqrt{I_{1}/{(I_{1}+I_{2,1})}}\leq-\sqrt{I_{1}/{(I_{1}+I_{2,2})}}=-\mathrm{Cov}(Z_{1},Z(I_{2,2}))=\mathrm{Cov}(Y_{1},Y_{2})$. Slepian’s inequality now gives for all $a_{1},a_{2}\in\mathbb{R}$ that $\mathbb{P}_{\delta}(X_{1}<a_{1},X_{2}\leq a_{2})\leq\mathbb{P}_{\delta}(Y_{1}<a_{1},Y_{2}\leq a_{2})$. By setting $a_{1}=z_{f}-\delta\sqrt{I_{1}}$ and $a_{2}=-z_{\alpha}+\delta\sqrt{I_{1}+I_{2,1}}$ and using $\delta>0$ which implies $a_{2}<-z_{\alpha}+\delta\sqrt{I_{1}+I_{2,2}}$, we obtain
@@
$<\mathbb{P}_{\delta}(Y_{1}<a_{1},Y_{2}\leq-z_{\alpha}+\delta\sqrt{I_{1}+I_{2,2}})=\mathbb{P}_{\delta}(Z_{1}<z_{f},Z(I_{2,2})\geq z_{\alpha})\penalty 10000\ .$
@@ C.2 Without binding futility stop
In this Section, it is assumed as in Section 4 that at the time of the planning of a fast-track procedure by the applicants, a successful conditional registration is a requirement for permanent registration. Thus, we assume that $\mathbb{P}_{\delta}(Z_{1}\geq z_{f},\text{\penalty 10000\ reject\penalty 10000\ }H_{0})=1-\beta$ is striven at an a priori assumed or true effect $\delta=\xi\cdot\delta_{\text{rel}}$ where $\xi>\xi_{\min}$. We assume that the designs and informations are chosen such that this success probability is achieved, as in Section 4. In contrast to Section 4, the conditional error functions are not chosen in a way that there is a binding futility stop after unsuccessful conditional registration. Instead, in this section, we deal with this case like an unbinding futility stop and consider $A_{I,-\infty}$ and $A_{F,-\infty}$. These are somewhat conservative designs with somewhat larger sample sizes compared to $A_{I,z_{f}}$ and $A_{F,z_{f}}$. However, such designs provide the opportunity to continue with a second stage for the permanent registration whenever there are good reasons for and sufficient resources have been achieved.
@@
Figure 11: Minimum needed information for the second stage to achieve overall power of $1-\beta=0.8$ for the common non-adaptive design and the adaptive designs with conditional error functions $A_{I,-\infty}$ (inverse normal method) and $A_{F,-\infty}$ (Fisher’s product test).
@@ Appendix D Additional Plots for Section 5
The plots in this section complete the discussion in Section 5. Four further plots for the effect assumption $\xi=1.25$ are presented. Additionally, all efficiency parameters considered in Section 5 for $\xi=1.25$ are calculated under the assumption $\xi=1.43=\xi_{\min}$.
Figure 16: Constant second stage information after non-successful conditional registration (left) and minimum second stage information after successful conditional registration (right) for $\xi=1.25$.